OpenStudy (anonymous):

suppose that a small pump can empty a swimming pool in 90 hours and that a large pump can empty a pool in 60 hours. Working together how long will it take them to fill the order?

OpenStudy (anonymous):

1/ 60 + 1/90 = (2 + 3)/ 180 = 5/180 => Total time is 180/5 = 36 hrs to empty the swimming pool.